4. What is the minimum deposit you would need to reduce monthly repayments to $150 on the best value loan?


Deposit: $10,000

Rate: 5.95%

Time: 5 years

Compounding Periods: 12


1
Expert's answer
2021-03-30T13:42:51-0400

This is an annuity:

"FV=A\frac{(1+r)^n-1}{r}"

"FV=150\frac{(1+0.0595)^{12}-1}{0.0595}=2523.12"